# Zoom-IQA: Image Quality Assessment with Reliable Region-Aware Reasoning

### v1

@@ Abstract
Image Quality Assessment (IQA) is a long-standing problem in computer vision. Previous methods typically focus on predicting numerical scores without explanation or provide low-level descriptions lacking precise scores. Recent reasoning-based vision language models (VLMs) have shown strong potential for IQA, enabling joint generation of quality descriptions and scores. However, we notice that existing VLM-based IQA methods tend to exhibit unreliable reasoning due to their limited capability of integrating visual and textual cues. In this work, we introduce Zoom-IQA, a VLM-based IQA model to explicitly emulate key cognitive behaviors: uncertainty awareness, region reasoning, and iterative refinement. Specifically, we present a two-stage training pipeline: 1) supervised fine-tuning (SFT) on our Grounded-Rationale-IQA (GR-IQA) dataset to teach the model to ground its assessments in key regions; and 2) reinforcement learning (RL) for dynamic policy exploration, primarily stabilized by our KL-Coverage regularizer to prevent reasoning and scoring diversity collapse, and supported by a Progressive Re-sampling Strategy to mitigate annotation bias. Extensive experiments show that Zoom-IQA achieves improved robustness, explainability, and generalization. The application to downstream tasks, such as image restoration, further demonstrates the effectiveness of Zoom-IQA.
Figure 1: (Upper) Current IQA methods are non-interactive, leading to inferior assessments. They either spot only partial flaws (e.g., slightly overexposed or slightly blurred) or make factually incorrect claims (clear and well-lit), resulting in erroneous judgments. Our Zoom-IQA uses interactive, region-aware reasoning: it first hypothesizes flaws (green text), then grounds them by cropping (orange text), and finally verifies the degradation (blue text). This hypothesize-and-verify loop provides a complete and accurate assessment. (Lower) Our model’s reasoning outputs also benefit downstream tasks, such as text-guided image restoration with SUPIR [yu2024scaling]. Our prompt enables a far superior restoration compared to those guided by other IQA methods or SUPIR’s default VLM, LLaVA-1.5-13b [liu2023visual].
11footnotetext: Corresponding author (shangchenzhou@gmail.com)
@@ 1 Introduction
Image Quality Assessment (IQA) is a fundamental task in computer vision, aiming to evaluate the perceptual quality of images in alignment with human perception. Its importance has grown rapidly as IQA models increasingly serve as critical perceptual reward signals for improving modern algorithms. Specifically, IQA serves as a component in frameworks like Reinforcement Learning (RL) from Human Feedback (RLHF) [rombach2022high, dhariwal2021diffusion, wang2025unified, he2024videoscore] to align outputs with human preferences. Similarly, in image restoration, IQA scores are used as differentiable rewards [sun2024rfsr] or in Direct Preference Optimization (DPO) [wu2025dp, cai2025dspo] to guide models toward perceptually superior results.
@@
The emergence of vision language models (VLMs) like CLIP [radford2021learning] opened a promising direction for IQA [wang2022exploring], further advanced by large-scale VLMs [touvron2023llama, liu2023visual, bai2025qwen2] that leverage broad knowledge to better align with human perception. Existing VLM-based IQA methods fall into two categories: (1) Score-based methods (e.g., Q-Align [wu2023q] and DeQA-Score [you2025teaching]), which emphasize accurate scores but lack the ability to provide textual descriptions; and (2) Description-based methods (e.g., DepictQA series [you2024depicting, you2024descriptive]), which offer detailed explanations but rely on SFT data with descriptions generated from ground-truth synthetic distortions. To bridge this gap, recent methods like Q-Insight [li2025q] and VisualQualityR1 [wu2025visualquality] introduce RL to unify quality scoring and textual reasoning using only score labels as rewards.
@@
Despite impressive capabilities, these VLM-based IQA methods remain non-interactive. They generate responses in a single pass without any mechanism for iterative visual refinement or correction. As highlighted in complex visual tasks (e.g., object detection [liu2025visual, shen2025vlm] and visual question answering (VQA) [yu2025perception, meng2025mm]), the lack of intermediate visual grounding may lead to unreliable responses and reasoning, especially under complex scenarios. Similarly, the nature of IQA requires subjects to interpret complex images by “zooming in” on key regions. The importance of this behavior is highlighted by the DiffIQA dataset [chen2025toward], which provided annotators with a zoom-in feature to inspect details. Such a region-aware interaction plays a helpful role in understanding the quality of an image, but is absent in existing VLM-based IQA models, confining their reasoning to the text domain and limiting effective use of visual information, as shown in Fig. 1.
@@
Enabling a VLM-based IQA method to dynamically “crop and zoom” for iterative region-aware assessment faces two core challenges. (1) Region-aware Learning. The model must learn where to focus and how to transform regions (e.g., crop, zoom) based on its own partial textual deliberations, similar to the grounding in VQA [zheng2025deepeyes, shao2024visual]. However, grounding IQA is more challenging than VQA. In VQA, grounding is often explicit, i.e., an answer can be tied to discrete, localizable semantic objects, for which extensive annotations are available [zheng2025deepeyes, shao2024visual]. In contrast, an IQA score is a holistic judgment aggregated from numerous, complex factors. The core difficulty is the lack of supervision specifying which regions a human prioritized to arrive at their final score. While recent IQA grounding datasets [chen2024q, chen2024grounding] provide static distortion masks, these methods fail to reveal the criticality of those regions to the overall human assessment or capture the dynamic reasoning path that led to the final judgment. (2) Self-Guided Reasoning Policy. The model must learn a dynamic policy on when to trigger detailed visual inspection rather than relying on random exploration or preset rules for zooming. This requires an unsupervised iterative cognitive process, i.e., performing a holistic assessment, identifying its own uncertainty about a specific region, and only then deciding to “zoom in” for refinement.
@@
To bridge such gaps, we make two primary contributions. First, we introduce Grounded-Rationale-IQA (GR-IQA), a fine-grained dataset curated to facilitate the development of interleaved text-image Chain-of-Thought (CoT) reasoning. Directly harnessing advanced VLMs such as Gemini [comanici2025gemini] to label IQA data with reasoning and scores can suffer from misalignment between visual inputs and reasoning outputs due to hallucination [leng2024mitigating, liu2024paying]. Our GR-IQA is designed to avoid such hallucination by providing rationales that are verifiably grounded in visual regions. Specifically, our curation pipeline introduces two key modules: (1) Visual Reliance Filtering (VRF), which enforces grounding by measuring the generative output shift (with vs. without the image); and (2) Hint-Augmented Consistency Filtering (HACF), which filters hallucination-like descriptions at the sentence level to maintain fluency. Second, we propose Zoom-IQA (Zoomable Region Reasoning for Reliable Image Quality Assessment), a novel framework designed to enhance the reasoning reliability of IQA with region awareness. Zoom-IQA is trained in two stages: it first learns formatted grounding (how to “zoom”) via supervised fine-tuning (SFT) on our GR-IQA dataset, and then learns a dynamic policy (when to “zoom”) via reinforcement learning (RL). Such learned policy allows Zoom-IQA to operate iteratively, moving beyond “single-pass” methods to identify uncertainty and refine its assessment, achieving truly interactive visual reasoning. To stabilize the training process, we further propose the KL-Coverage regularizer, designed to prevent a collapse in reasoning path diversity, which often leads to a severe “mode collapse” in predicted scores. A Progressive Re-sampling Strategy is also developed to mitigate bias from imbalanced annotations.
@@ 2 Related Work
Image Quality Assessment. Previous IQA works are broadly divided into full-reference (FR) and no-reference (NR) approaches, based on the availability of a pristine reference image. As our work does not require a reference, we focus on the more challenging NR-IQA task. Conventional NR-IQA methods [ma2017learning, mittal2012no, mittal2012making, moorthy2010two, moorthy2011blind] relied on hand-crafted, degradation-aware features to predict the final quality score. Subsequent deep learning-based models [kang2014convolutional, bosse2017deep, su2020blindly, ke2021musiq, cheon2021perceptual, sun2022graphiqa, liu2017rankiqa, pan2018blind, zhu2020metaiqa] replaced this pipeline, directly predicting quality scores using end-to-end trainable neural networks. Nonetheless, these models often suffer from significant performance degradation on out-of-distribution (OOD) data, limiting their practical usage and generalizability.
@@
Vision Language Models in Image Quality Assessment. Vision language models (VLMs) [radford2021learning, touvron2023llama, liu2023visual, bai2025qwen2] have been extensively studied in image quality assessment (IQA), leveraging their powerful cross-modal understanding and strong generalization capabilities. These works typically focus on one of two objectives: providing numerical quality scores [wang2022exploring, wu2023q, you2025teaching, zhu2024adaptive, liu2024dog] or generating visual quality descriptions [wu2024q, you2024depicting, you2024descriptive, chen2024grounding]. Specifically, CLIP-IQA proposes to harness CLIP [radford2021learning] for image quality assessment from multiple aspects. DOG-IQA [liu2024dog] attempts to mimic the human evaluation process (e.g., zooming in to evaluate specific areas). However, these models lack the necessary reasoning capabilities and cannot dynamically decide which regions to inspect. A common way to alleviate such limitations is via pre-processing, i.e., using a pre-trained segmentation model to crop sub-images and then computing the final score as a weighted average of these crops. Recent works like Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality], and Q-Ponder [cai2025q] propose to employ reinforcement learning (RL) to leverage the reasoning capabilities of VLMs, enabling image quality rating as well as textual justifications. However, the reasoning chains generated by these methods remain purely textual. They lack dynamic interaction with the image (e.g., cropping or zooming) to evaluate specific regions—a process crucial to human assessment. Consequently, visual evidence is insufficiently explored, and the models fail to ground their textual reasoning in verifiable image regions, limiting both the reliability and interpretability of their outputs.
@@
Vision Language Models with Multimodal Reasoning. Recent advances in enhancing the reasoning capabilities of VLMs have significantly improved the performance of VLMs on challenging tasks, such as mathematical problem solving [meng2025mm, huang2025vision, zhang2025r1], VQA [yu2025perception, meng2025mm], and object detection [liu2025visual, shen2025vlm, yu2025perception]. However, these models typically generate reasoning chains composed solely of natural language. This text-only reasoning can be opaque and often lacks sufficient grounding in the visual input’s fine-grained details. To address this, recent works in VQA [zheng2025deepeyes, zhang2025chain, su2025pixel, jiang2025vlm] propose to integrate evidence regions into the reasoning process. By equipping VLMs with capabilities like iterative zoom-in and region-of-interest selection, these methods demonstrated boosted performance. Improved interpretability is further gained via visual-linguistic interactive reasoning. However, these VQA methods generally rely on heavily annotated evidence regions (e.g., bounding boxes) for training. Such fine-grained regional labels and corresponding reasoning trajectories are critically expensive and lacking in the IQA domain.
@@ 3 Methodology
Our method, Zoom-IQA, is trained via a two-stage pipeline (illustrated in Fig. 2): 1) Supervised Fine-Tuning for Grounded Quality Rationale Learning: We first leverage our GR-IQA dataset to teach the VLM the foundational “how-to” skills: grounding textual rationales in visual regions and executing the “zoom” action (Sec. 3.1). 2) Reinforcement Learning for Self-Guided Exploration: To enable iterative refinement without exhaustive supervision, we employ Reinforcement Learning to derive a dynamic policy that optimizes the deployment of these skills. Specifically, we adopt Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] as our optimization framework, facilitating stable and efficient policy exploration (Sec. 3.2).
@@ 3.1 Grounded Quality Rationale Learning
Data Generation. We prompt the closed-source VLM, Gemini-2.5-pro [comanici2025gemini], on KonIQ dataset [hosu2020koniq] images using a structured prompt. This compels the VLM to generate a two-part response: a textual rationale (within <think>) and a JSON action (within <answer>). The textual rationale is strictly constrained to a four-part format: (1) a holistic Image Quality Summary; (2) Directions for Improvement; (3) Issues to Avoid; and (4) a Decision & Rationale where the model must decide if the image is an “easy case” (requiring a “final” tool) or a “hard case” (requiring a “crop” tool) and justify its choice. The <answer> block then contains the chosen “tool” (“final” or “crop”), the rating, and a conditional “bbox”. This structured process forces the VLM to link its score to regional evidence and to perform self-assessment on its own uncertainty (i.e., “zoom” or not). This raw output forms the data for our GR-IQA dataset, which is then passed to our filtering modules (VRF and HACF).
Visual Reliance Filtering (VRF). Since VLMs inherit strong language modeling capabilities from LLMs, they may over-rely on textual co-occurrence patterns, leading to hallucinated outputs [liu2024paying, leng2024mitigating]. To address this, we propose Visual Reliance Filtering (VRF). This module filters low-reliance samples by comparing the VLM’s output in two distinct scenarios: 1) conditioned on both the image $I$ and the textual rationale $R_{A}$, and 2) conditioned only on the rationale $R_{A}$. If the outputs are too similar, we discard the sample, as this indicates the visual input $I$ was not essential and the answer could be reproduced from the text alone.
Hint-Augmented Consistency Filtering (HACF). While VRF ensures the final <answer> is grounded, the textual rationale $R$ (the <think> block) may still contain unfaithful statements. To filter these low-veracity samples, we employ a powerful LLM, Qwen-2.5-32b [qwen2.5], denoted $\text{LLM}_{\text{Rater}}$, to perform a holistic assessment of the entire rationale $R$. Crucially, to aid this judgment, the $\text{LLM}_{\text{Rater}}$ is provided not only with the full rationale $R$ and the image $I$, but also with a set of pre-computed, low-level hints $H$ (e.g., global brightness, sharpness, color metrics). The rater then outputs a single, binary veracity decision $D_{\text{HACF}}$ (i.e., “Pass” or “Fail”) for the entire sample: $D_{\text{HACF}}=\text{LLM}_{\text{Rater}}(R,I,H)$ We only retain samples with $D_{\text{HACF}}=\text{``Pass"}$, ensuring that only data with high-veracity reasoning is kept for training.

### v2

@@ Abstract
Image Quality Assessment (IQA) is a long-standing problem in computer vision. Previous methods typically focus on predicting numerical scores without explanation or providing low-level descriptions lacking precise scores. Recent reasoning-based vision language models (VLMs) have shown strong potential for IQA by jointly generating quality descriptions and scores. However, existing VLM-based IQA methods often suffer from unreliable reasoning due to their limited capability of integrating visual and textual cues. In this work, we introduce Zoom-IQA, a VLM-based IQA model to explicitly emulate key cognitive behaviors: uncertainty awareness, region reasoning, and iterative refinement. Specifically, we present a two-stage training pipeline: 1) supervised fine-tuning (SFT) on our Grounded-Rationale-IQA (GR-IQA) dataset to teach the model to ground its assessments in key regions, and 2) reinforcement learning (RL) for dynamic policy exploration, stabilized by our KL-Coverage regularizer to prevent reasoning and scoring diversity collapse, with a Progressive Re-sampling Strategy for mitigating annotation bias. Extensive experiments show that Zoom-IQA achieves improved robustness, explainability, and generalization. The application to downstream tasks, such as image restoration, further demonstrates the effectiveness of Zoom-IQA.
Figure 1: (Upper) Current IQA methods are non-interactive, leading to inferior assessments. They either spot only partial flaws (e.g., slightly overexposed or slightly blurred) or make factually incorrect claims (clear and well-lit), resulting in erroneous judgments. Our Zoom-IQA uses interactive, region-aware reasoning: it first hypothesizes flaws (green text), then grounds them by cropping (orange text), and finally verifies the degradation (blue text). This hypothesize-and-verify loop provides a complete and accurate assessment. (Lower) Our model’s reasoning outputs also benefit downstream tasks, such as text-guided image restoration with SUPIR [75]. Our prompt enables a far superior restoration compared to those guided by other IQA methods or SUPIR’s default VLM, LLaVA-1.5-13b [32].
11footnotetext: Corresponding author
@@ 1 Introduction
Image Quality Assessment (IQA) is a fundamental task in computer vision, aiming to evaluate the perceptual quality of images in alignment with human perception. Its importance has grown rapidly as IQA models increasingly serve as critical perceptual reward signals for improving modern algorithms. Specifically, IQA serves as a component in frameworks like Reinforcement Learning (RL) from Human Feedback (RLHF) [46, 14, 63, 17] to align outputs with human preferences. Similarly, in image restoration, IQA scores are used as differentiable rewards [55] or in Direct Preference Optimization (DPO) [66, 4] to guide models toward perceptually superior results.
@@
The emergence of vision language models (VLMs) like CLIP [44] opened a promising direction for IQA [59], further advanced by large-scale VLMs [58, 32, 2] that leverage broad knowledge to better align with human perception. Existing VLM-based IQA methods fall into two categories: (1) Score-based methods (e.g., Q-Align [65] and DeQA-Score [71]), which emphasize accurate scores but lack the ability to provide textual descriptions; and (2) Description-based methods (e.g., DepictQA series [73, 72]), which offer detailed explanations but rely on SFT data with descriptions generated from ground-truth synthetic distortions. To bridge this gap, recent methods like Q-Insight [29] and VisualQualityR1 [67] introduce RL to unify quality scoring and textual reasoning using only score labels as rewards.
@@
Despite impressive capabilities, these VLM-based IQA methods remain non-interactive. They generate responses in a single pass without any mechanism for iterative visual refinement or correction. As highlighted in complex visual tasks (e.g., object detection [36, 50] and visual question answering (VQA) [74, 38]), the lack of intermediate visual grounding may lead to unreliable responses and reasoning, especially under complex scenarios. Similarly, the nature of IQA requires subjects to interpret complex images by “zooming in” on key regions. The importance of this behavior is highlighted by the DiffIQA dataset [8], which provided annotators with a zoom-in feature to inspect details. Such a region-aware interaction plays a helpful role in understanding the quality of an image, but is absent in existing VLM-based IQA models, confining their reasoning to the text domain and limiting effective use of visual information, as shown in Fig. 1.
@@
Enabling a VLM-based IQA method to dynamically “crop and zoom” for iterative region-aware assessment faces two core challenges. (1) Region-aware Learning. The model must learn where to focus and how to transform regions (e.g., crop, zoom) based on its own partial textual deliberations, similar to the grounding in VQA [79, 48]. However, grounding IQA is more challenging than VQA. In VQA, grounding is often explicit, i.e., an answer can be tied to discrete, localizable semantic objects, for which extensive annotations are available [79, 48]. In contrast, an IQA score is a holistic judgment aggregated from numerous, complex factors. The core difficulty is the lack of supervision specifying which regions a human prioritized to arrive at their final score. While recent IQA grounding datasets [6, 9] provide static distortion masks, these methods fail to reveal the criticality of those regions to the overall human assessment or capture the dynamic reasoning path that led to the final judgment. (2) Self-Guided Reasoning Policy. The model must learn a dynamic policy on when to trigger detailed visual inspection rather than relying on random exploration or preset rules for zooming. This requires an unsupervised iterative cognitive process, i.e., performing a holistic assessment, identifying its own uncertainty about a specific region, and only then deciding to “zoom in” for refinement.
@@
To bridge such gaps, we make two primary contributions. First, we introduce Grounded-Rationale-IQA (GR-IQA), a fine-grained dataset curated to facilitate the development of interleaved text-image Chain-of-Thought (CoT) reasoning. Directly harnessing advanced VLMs such as Gemini [12] to label IQA data with reasoning and scores can suffer from misalignment between visual inputs and reasoning outputs due to hallucination [27, 34]. Our GR-IQA is designed to avoid such hallucination by providing rationales that are verifiably grounded in visual regions. Specifically, our curation pipeline introduces two key modules: (1) Visual Reliance Filtering (VRF), which enforces grounding by measuring the generative output shift (with vs. without the image); and (2) Hint-Augmented Consistency Filtering (HACF), which filters hallucination-like descriptions at the sentence level to maintain fluency. Second, we propose Zoom-IQA (Zoomable Region Reasoning for Reliable Image Quality Assessment), a novel framework designed to enhance the reasoning reliability of IQA with region awareness. Zoom-IQA is trained in two stages: it first learns formatted grounding (how to “zoom”) via supervised fine-tuning (SFT) on our GR-IQA dataset, and then learns a dynamic policy (when to “zoom”) via reinforcement learning (RL). Such learned policy allows Zoom-IQA to operate iteratively, moving beyond “single-pass” methods to identify uncertainty and refine its assessment, achieving truly interactive visual reasoning. To stabilize the training process, we further propose the KL-Coverage regularizer, designed to prevent a collapse in reasoning path diversity, which often leads to a severe “mode collapse” in predicted scores. A Progressive Re-sampling Strategy is also developed to mitigate bias from imbalanced annotations.
@@ 2 Related Work
Image Quality Assessment. Previous IQA works are broadly divided into full-reference (FR) and no-reference (NR) approaches, based on the availability of a pristine reference image. As our work does not require a reference, we focus on the more challenging NR-IQA task. Conventional NR-IQA methods [37, 39, 40, 41, 42] relied on hand-crafted, degradation-aware features to predict the final quality score. Subsequent deep learning-based models [24, 3, 53, 25, 11, 54, 35, 43, 80] replaced this pipeline, directly predicting quality scores using end-to-end trainable neural networks. Nonetheless, these models often suffer from significant performance degradation on out-of-distribution (OOD) data, limiting their practical usage and generalizability.
@@
Vision Language Models in Image Quality Assessment. Vision language models (VLMs) [44, 58, 32, 2] have been extensively studied in image quality assessment (IQA), leveraging their powerful cross-modal understanding and strong generalization capabilities. These works typically focus on one of two objectives: providing numerical quality scores [59, 65, 71, 81, 33] or generating visual quality descriptions [64, 73, 72, 9]. Specifically, CLIP-IQA proposes to harness CLIP [44] for image quality assessment from multiple aspects. DOG-IQA [33] attempts to mimic the human evaluation process (e.g., zooming in to evaluate specific areas). However, these models lack the necessary reasoning capabilities and cannot dynamically decide which regions to inspect. A common way to alleviate such limitations is via pre-processing, i.e., using a pre-trained segmentation model to crop sub-images and then computing the final score as a weighted average of these crops. Recent works like Q-Insight [29], VisualQuality-R1 [67], and Q-Ponder [5] propose to employ reinforcement learning (RL) to leverage the reasoning capabilities of VLMs, enabling image quality rating as well as textual justifications. However, the reasoning chains generated by these methods remain purely textual. They lack dynamic interaction with the image (e.g., cropping or zooming) to evaluate specific regions—a process crucial to human assessment. Consequently, visual evidence is insufficiently explored, and the models fail to ground their textual reasoning in verifiable image regions, limiting both the reliability and interpretability of their outputs.
@@
Vision Language Models with Multimodal Reasoning. Recent advances in enhancing the reasoning capabilities of VLMs have significantly improved the performance of VLMs on challenging tasks, such as mathematical problem solving [38, 19, 76], VQA [74, 38], and object detection [36, 50, 74]. However, these models typically generate reasoning chains composed solely of natural language. This text-only reasoning can be opaque and often lacks sufficient grounding in the visual input’s fine-grained details. To address this, recent works in VQA [79, 78, 51, 21] propose to integrate evidence regions into the reasoning process. By equipping VLMs with capabilities like iterative zoom-in and region-of-interest selection, these methods demonstrated boosted performance. Improved interpretability is further gained via visual-linguistic interactive reasoning. However, these VQA methods generally rely on heavily annotated evidence regions (e.g., bounding boxes) for training. Such fine-grained regional labels and corresponding reasoning trajectories are critically expensive and lacking in the IQA domain.
@@ 3 Methodology
Our method, Zoom-IQA, is trained via a two-stage pipeline (illustrated in Fig. 2): 1) Supervised Fine-Tuning for Grounded Quality Rationale Learning: We first leverage our GR-IQA dataset to teach the VLM the foundational “how-to” skills: grounding textual rationales in visual regions and executing the “zoom” action (Sec. 3.1). 2) Reinforcement Learning for Self-Guided Exploration: To enable iterative refinement without exhaustive supervision, we employ Reinforcement Learning to derive a dynamic policy that optimizes the deployment of these skills. Specifically, we adopt Group Relative Policy Optimization (GRPO) [49] as our optimization framework, facilitating stable and efficient policy exploration (Sec. 3.2).
@@ 3.1 Grounded Quality Rationale Learning
Data Generation. We prompt the closed-source VLM, Gemini-2.5-pro [12], on KonIQ dataset [18] images using a structured prompt. This comple the VLM to generate a two-part response: a textual rationale (within <think>) and a JSON action (within <answer>). The textual rationale is strictly constrained to a four-part format: (1) a holistic Image Quality Summary; (2) Directions for Improvement; (3) Issues to Avoid; and (4) a Decision & Rationale where the model must decide if the image is an “easy case” (requiring a “final” tool) or a “hard case” (requiring a “crop” tool) and justify its choice. The <answer> block then contains the chosen “tool” (“final” or “crop”), the rating, and a conditional “bbox”. This structured process encourages the VLM to link its score to regional evidence and to self-assess uncertainty (i.e., “zoom” or not). This raw output forms the data for our GR-IQA dataset, which is then passed to our filtering modules (VRF and HACF).
Visual Reliance Filtering (VRF). Since VLMs inherit strong language modeling capabilities from LLMs, they may over-rely on textual co-occurrence patterns, leading to hallucinated outputs [34, 27]. To address this, we propose Visual Reliance Filtering (VRF). This module filters low-reliance samples by comparing the VLM’s output in two distinct scenarios: 1) conditioned on both the image $I$ and the textual rationale $R_{A}$, and 2) conditioned only on the rationale $R_{A}$. If the outputs are too similar, we discard the sample, as this indicates the visual input $I$ was not essential and the answer could be reproduced from the text alone.
Hint-Augmented Consistency Filtering (HACF). While VRF ensures the final <answer> is grounded, the textual rationale $R$ (the <think> block) may still contain unfaithful statements. To filter these low-veracity samples, we employ a powerful LLM, Qwen-2.5-32b [69], denoted $\text{LLM}_{\text{Rater}}$, to perform a holistic assessment of the entire rationale $R$. Crucially, to aid this judgment, the $\text{LLM}_{\text{Rater}}$ is provided not only with the full rationale $R$ and the image $I$, but also with a set of pre-computed, low-level hints $H$ (e.g., global brightness, sharpness, color metrics). The rater then outputs a single, binary veracity decision $D_{\text{HACF}}$ (i.e., “Pass” or “Fail”) for the entire sample: $D_{\text{HACF}}=\text{LLM}_{\text{Rater}}(R,I,H)$ We only retain samples with $D_{\text{HACF}}=\text{``Pass"}$, ensuring that only data with high-veracity reasoning is kept for training.